# The Optimal Strategy for Playing Lucky 13

### v1

@@ 10 Conclusion and Future Work
We have presented models for the optimal strategy of Lucky 13 with 2 and 3 question categories based on two different utility functions. We have also presented two case studies of actual games and used Monte Carlo simulation to estimate the number of questions an average player may answer correctly. Additionally, we explain how this model could be extended to allow contestants to select individual probabilities of giving a correct answer for each of the 13 questions. This research could be extended by conducting an empirical study to determine if contestant decisions in a real game came close to our optimal strategy while playing. This idea is subject to ongoing research. It would also be interesting to model the amount of money the producers offer a contestant to quit. However, the amount of data available was too sparse to come up with a realistic model.

### v2

@@ 11 Conclusion and Future Work
We have presented models for the optimal strategy of Lucky 13 with 2 and 3 question categories based on two different utility functions. We have also presented two case studies of actual games and used Monte Carlo simulation to estimate the number of questions an average player may answer correctly. Additionally, we explain how this model could be extended to allow contestants to provide individual probabilities of giving a correct answer for each of the Unsure questions or all 13 questions. This research could be extended by conducting an empirical study to determine if contestant decisions in a real game came close to our optimal strategy while playing. Furthermore, a Monte Carlo simulation could be conducted to estimate the probabilities (and their standard errors) for each Lucky Number under the extended three category model given the contestant data and an appropriate prior distribution. Both of these ideas are subject to ongoing research. It would also be interesting to model the amount of money the producers offer a contestant to quit. However, the amount of data available was too sparse to come up with a realistic model.
12 Statements and Declarations
12.1 Author Contribution
The manuscript was prepared by S.B. and all authors reviewed the manuscript and results.
12.2 Funding
The authors have no relevant financial or non-financial interests to disclose.
12.3 Ethics Approval
Not applicable. This research does not qualify as human subject research.